The length of a rectangle is 11 more than twice its width. The perimeter of the rectangle is 118 feet. Find the dimensions of the rectangle
Answer by hamsanash1981@gmail.com(151)

let the width be x, then length = 2x+11
perimeter = 118
2(2X+11+X) =118
3X+11 = 59
3x = 48
x = 16
width = 16
length = 2(16)+11 = 43
the dimension of the rectangle = 43*16